OpenStudy (anonymous):

A sofa and a chair cost $850 for both. If when sold individually the sofa costs $100 more than twice the price of the chair, what is the price of each piece of furniture?

OpenStudy (anonymous):

I am thinking $600 for the sofa and $250 for the chair.

OpenStudy (anonymous):

Alright so you know that the total cost will end up being $850. So to set up this question, key info is given such as the sofa cost 100 more than twice the price of the chair. So from this we need to create two equations. Before doing so: Let x represent the price of the sofa. Let y represent the price of a chair. 1) Together they cost 850 so one equation would be: " x + y= 850 " 2) Tip 2: The sofa cost 100 more than twice the price of the chair. " x= 2y + 100" From here we can simply use the substitution method to find the value of x and y. Plug 2) into 1) and we get: 2y + 100 + y = 850 3y+100 = 850 3y = 750 y = 250 Now sub this newly found value back into 1) x + y = 850 x +250 = 850 x = 600

OpenStudy (anonymous):

Hope this helped you understand the question :D!

OpenStudy (anonymous):

It did, because I just picked what looked resonable.